# pMixFed: Efficient Personalized Federated Learning through Adaptive Layer-Wise Mixup

### v1

@@ Abstract
Traditional Federated Learning (FL) methods encounter significant challenges when dealing with heterogeneous data and providing personalized solutions for non-IID scenarios. Personalized Federated Learning (PFL) approaches aim to address these issues by balancing generalization and personalization, often through parameter decoupling or partial models that freeze some neural network layers for personalization while aggregating other layers globally. However, existing methods still face challenges of global-local model discrepancy, client drift, and catastrophic forgetting, which degrade model accuracy. To overcome these limitations, we propose pMixFed, a dynamic, layer-wise PFL approach that integrates mixup between shared global and personalized local models. Our method introduces an adaptive strategy for partitioning between personalized and shared layers, a gradual transition of personalization degree to enhance local client adaptation, improved generalization across clients, and a novel aggregation mechanism to mitigate catastrophic forgetting. Extensive experiments demonstrate that pMixFed outperforms state-of-the-art PFL methods, showing faster model training, increased robustness, and improved handling of data heterogeneity under different heterogeneous settings. Our code is available for reproducing our results: https://github.com/YasMinSdt/pMixFed.
@@ 1 Introduction
The goal in federated learning (FL) [22] is to facilitate collaborative learning of several machine learning (ML) models in a decentralized scheme. FL requires addressing data privacy, catastrophic forgetting, and client drift problem  [^1] A phenomenon where the global model fails to serve as an accurate representation because local models gradually drift apart due to high data heterogeneity. [38, 15, 43, 31, 36]. Existing FL methods cannot address all these challenges with non-IID (non-Independent and Identically Distributed) data. For instance, although “FedAvg”  [32] demonstrates strong generalization performance, it fails to provide personalized solutions for a cohort of clients with non-IID datasets. Hence, the global model, or one “average client” in “FedAvg”, may not adequately represent all individual local models in non-IID settings due to client-drift [53]. Personalized FL (PFL) methods handle data heterogeneity by considering both generalization and personalization during the training stage. Since there is a trade-off between generalization and personalization in heterogeneous environments, PFL methods leverage heterogeneity and diversity as advantages rather than adversities [35, 47]. A group of PFL approaches train personalized local models on each device while collaborating toward a shared global model. Partial PFL, also known as parameter decoupling, involves using a partial model sharing, where only a subset of the model is shared while other parameters remain “frozen” to balance generalization and personalization until the subsequent round of local training.
@@
While partial PFL methods are effective in mitigating catastrophic forgetting, strengthening privacy, and reducing computation and communication overhead [34, 45], there are still some unaddressed challenges. First, the question of when, where, and how to optimally partition the full model? remains unresolved. Recent studies [34, 45] have shown that there is no “one-size-fits-all” solution; the best or optimal partitioning strategy depends on factors such as task type (e.g., next word prediction or speech recognition) and local model architecture. An improper partitioning choice can lead to issues such as underfitting, overfitting, increased bias, and catastrophic forgetting. Some studies [29] suggest that personalized layers should reside in the base layers, while others [9, 4] argue that the base layers contain more generalized information and should be shared. Further, the use of a fixed partitioning strategy across all communication rounds for heterogeneous clients can limit the efficacy of collaborative learning. For instance, if the performance of the client suddenly drops due to new incoming data, the partitioning strategy should be changed because the client requires more frozen layers. Another issue is catastrophic forgetting of the previously shared global knowledge after only a few rounds of local training because the shared global model can be completely overwritten by local updates leading to generalization degradation [31, 41, 15, 54]. Most importantly, partial models may experience slower convergence compared to full model personalization, as frozen local model updates can diverge in an opposite direction from the globally-shared model. Since the generalized and personalized models are trained on non-IID datasets, there might also be a domain shift, leading to model discrepancy as depicted in Figure 1. These discrepancies arise from variations in local and global objective functions, differences in initialization, and asynchronous updates [56, 24]. As a result, merging the shared and the personalized layers can disrupt information flow within the network, impede the learning process, and lead to a slower convergence rate or accuracy drop in partial PFL models such as FedAlt and FedSim [34] [^2] More details on this is discussed in section 4.. Further, while partial PFL techniques contribute to an overall improved training accuracy, they can reduce the test accuracy on some devices, particularly in devices with limited samples, leading to variations in results in terms of the performance level [34]. Hence, there is a need for novel solutions to achieve the following goals in PFL:
@@
  • Gradual Personalization Transition: The degree of personalization should transition gradually across layers, as opposed to an ”all-or-nothing” approach that employs strict partitioning or hard splits within the model discussed in Figure 1. This ability allows nuanced adaptation for individual client needs due to heterogeneity.
@@
To achieve the above, we propose “pMixFed”, a layer-wise, dynamic PFL approach that integrates Mixup [61] between the shared global and personalized local models’ layers during both the broadcasting (global model sharing with local clients) and aggregation (aggregating distributed local models to update the global model) stages within a partial PFL framework. Our main contributions include:
  • We develop an online, dynamic interpolation method between local and global models using Mixup [57], effectively addressing data heterogeneity and scalable across varying cohort sizes, degrees of data heterogeneity, and diverse model sizes and architectures.
@@ 2 Related Work
PFL aims to adapt local models to the individual needs, preferences, and contexts of each participant. Since the inception of FL, various PFL approaches have been explored to address the challenges stemming from heterogeneity, categorized into three areas: 1) data-centric strategies 2) global model adaptation, and 3) local model personalization. The following provide a more detailed discussion of these three approaches:
Data-centric: These approaches in PFL aim to address data heterogeneity and class imbalances by manipulating statistical data properties, such as data size, distribution, and local data selection [47]. Examples of these techniques include data normalization, feature engineering, data augmentation, synthetic data generation, and client selection. Representative methods, such as Astraea [11], P2P k-SMOTE [50], FedMCCS [1], FedAug [19], [63], and FedHome [52], exemplify data-centric approaches. The shortcoming of these methods is that they often modify the natural distribution and statistical characteristics of federated data, potentially injecting bias or eliminating valuable, rare information. Additionally, many of these strategies depend on the availability of proxy data on the global server, which may conflict with the privacy regulations of some entities.
@@
Global Model Adaptation: In these approaches, a single global model is maintained on the server and subsequently adapted to individual local models in the following phase. The primary goals of these methods are: 1) learning a robust, generalized model and 2) enabling fast, efficient local model adaptation. Several techniques employ regularization terms to mitigate client drift and prevent model divergence during local updates e.g., FedProx [27], SCAFFOLD [21], FL-MOON [26], and FedCurv [42]. Some of these approaches such as PerFedAvg [12], pFedMe [46], and pFedHN [39] leverage Meta-learning, and transfer learning e.g., FedStag [55], and DPFed [58], yet they introduce specific challenges. Meta-learning methods can be computationally intensive, while regularization techniques may add computational overhead by incorporating additional terms in the objective function. Similarly, transfer learning algorithms can be inefficient in terms of communication overhead and often require a public dataset to enhance the global model on the server. A common limitation across most adaptation techniques is the need for uniform model architecture across all clients, requiring devices with varying computational capabilities to use the same model size.
Local Model Personalization: The limitations of existing PFL methods have led to the development of local model personalization approaches that focus on training customized models for each client. Some approaches utilize multi-task learning (MTL), a collaborative framework that facilitates information exchange across distinct tasks (e.g., FedRes [2], MOCHA [44], FedAMP [16], and Ditto [28]). Another category leverages knowledge distillation (KD) to support personalization when client-specific training objectives differ, as seen in FedDF [30], FedMD [25], FedGen [64], and FedGKT [14]. However, both MTL and KD can incur significant computational and communication overhead, limiting their scalability for large-scale FL deployments across resource-constrained devices.
@@
Partial PFL: These methods, also known as parameter decoupling, are the primary focus of this paper. These methods mitigate catastrophic forgetting by retaining shared components while freezing other layers for personalization. FedPer [4] introduced partial models in FL, sharing only initial layers with generalized information and reserving final layers for personalization. FedBABU [33] divides the network into a shared body and a frozen head with fully connected layers. Other frameworks, such as FURL [6] and LG-FedAvg [29], apply partial PFL by retaining private feature embeddings or using compact representation learning for high-level features, respectively. Two baseline methods in this paper are FedAlt [43] and FedSim [34]. FedAlt uses a stateless FL paradigm, reconstructing local models from the global model, while FedSim synchronously updates shared and local models with each iteration.
@@
Limitations: These methods face challenges such as model update discrepancies (as shown in Figure 1) and catastrophic forgetting, where shared layers may undergo significant changes after only a few rounds of local training, resulting in sudden accuracy drops. Additionally, users with high personalization accuracy may freeze more layers than cold-start users or unreliable participants, who should rely more heavily on the global model. These challenges have motivated our development of a dynamic, adaptive layer-wise approach to balance generalization and personalization across clients, allowing tuning at different communication rounds to accommodate varying performance conditions.

### v2

@@ Abstract
Traditional Federated Learning (FL) methods encounter significant challenges when dealing with heterogeneous data and providing personalized solutions for non-IID scenarios. Personalized Federated Learning (PFL) approaches aim to address these issues by balancing generalization and personalization, often through parameter decoupling or partial models that freeze some neural network layers for personalization while aggregating other layers globally. However, existing methods still face challenges of global-local model discrepancy, client drift, and catastrophic forgetting, which degrade model accuracy. To overcome these limitations, we propose pMixFed, a dynamic, layer-wise PFL approach that integrates mixup between shared global and personalized local models. Our method introduces an adaptive strategy for partitioning between personalized and shared layers, a gradual transition of personalization degree to enhance local client adaptation, improved generalization across clients, and a novel aggregation mechanism to mitigate catastrophic forgetting. Extensive experiments demonstrate that pMixFed outperforms state-of-the-art PFL methods, showing faster model training, increased robustness, and improved handling of data heterogeneity under different heterogeneous settings.
Keywords: Personalization, Federated Learning, Meta Learning
@@ 1 Introduction
AI has achieved remarkable success across computer vision, natural language processing, and healthcare tasks by leveraging large-scale datasets. However, in many real-world scenarios data is distributed across multiple devices or organizations, and privacy ma2020safeguarding; mothukuri2021federated; stansecure; zhu2024enhancing, regulatory cai2024esvfl; pati2024privacy; nananukul2024multi; yazdinejad2024robust, or communication constraints yao2018two; luping2019cmfl; paragliola2022definition prevent centralizing the data. This challenge increases heterogeneity in data distributions across clients which can degrade the performance of centralized AI methods. Moreover, user-specific preferences or contexts often require personalized models rather than a single global solution. These practical necessities motivate the need for collaborative learning methods that allow training models jointly across clients while respecting privacy, reducing communication costs, and adapting to heterogeneous clients.
@@
One goal of federated learning (FL) konevcny2016federated is to facilitate collaborative learning of several machine learning (ML) models in a decentralized scheme. FL requires addressing data privacy, catastrophic forgetting, and client drift problem  [^1] A phenomenon where the global model fails to serve as an accurate representation because local models gradually drift apart due to high data heterogeneity. rostami2018multi; huang2022learn; singhal2021federated; luo2023gradma; qu2022rethinking. Existing FL methods cannot address all these challenges with non-Independent and Identically Distributed (non-IID) data. For instance, although Federated Averaging (“FedAvg”) mcmahan2017communication demonstrates strong generalization performance, it fails to provide personalized solutions for a cohort of clients with non-IID datasets. Hence, the global model, or one “average client” in “FedAvg”, may not adequately represent all individual local models in non-IID settings due to client-drift xiao2020averaging. Personalized FL (PFL) methods handle data heterogeneity by considering both generalization and personalization during the training stage. Since there is a trade-off between generalization and personalization in heterogeneous environments, PFL methods leverage heterogeneity and diversity as advantages rather than adversities pye2021personalised; tan2022towards. A group of PFL approaches train personalized local models on each device while collaborating toward a shared global model. Partial PFL, also known as parameter decoupling, involves using a partial model sharing, where only a subset of the model is shared while other parameters remain “frozen” to balance generalization and personalization until the subsequent round of local training.
@@
While partial PFL methods are effective in mitigating catastrophic forgetting, strengthening privacy, and reducing computation and communication overhead pillutla2022federated; sun2023fedperfix, there are still some unaddressed challenges. First, the question of when, where, and how to optimally partition the full model? remains unresolved. Recent studies pillutla2022federated; sun2023fedperfix have shown that there is no “one-size-fits-all” solution; the best or optimal partitioning strategy depends on factors such as task type (e.g., next word prediction or speech recognition) and local model architecture. An improper partitioning choice can lead to issues such as underfitting, overfitting, increased bias, and catastrophic forgetting. Some studies liang2020think suggest that personalized layers should reside in the base layers, while others collins2021exploiting; arivazhagan2019federated argue that the base layers contain more generalized information and should be shared. Further, the use of a fixed partitioning strategy across all communication rounds for heterogeneous clients can limit the efficacy of collaborative learning. For instance, if the performance of the client suddenly drops due to new incoming data, the partitioning strategy should be changed because the client requires more frozen layers. Another issue is catastrophic forgetting of the previously shared global knowledge after only a few rounds of local training because the shared global model can be completely overwritten by local updates leading to generalization degradation luo2023gradma; shirvani2023comprehensive; huang2022learn; xu2022acceleration. Most importantly, partial models may experience slower convergence compared to full model personalization, as frozen local model updates can diverge in an opposite direction from the globally-shared model. Since the generalized and personalized models are trained on non-IID datasets, there might also be a domain shift, leading to model discrepancy as depicted in Figure 1. These discrepancies arise from variations in local and global objective functions, differences in initialization, and asynchronous updates yang2024fedas; lee2023partial. As a result, merging the shared and the personalized layers can disrupt information flow within the network, impede the learning process, and lead to a slower convergence rate or accuracy drop in partial PFL models such as FedAlt and FedSim pillutla2022federated [^2] More details on this is discussed in section 4.. Further, while partial PFL techniques contribute to an overall improved training accuracy, they can reduce the test accuracy on some devices, particularly in devices with limited samples, leading to variations in results in terms of the performance level pillutla2022federated. Hence, there is a need for novel solutions to achieve the following goals in PFL:
@@
  • Gradual Personalization Transition: The degree of personalization should transition gradually across layers, as opposed to an “all-or-nothing” approach that employs strict partitioning or hard splits within the model discussed in Figure 1. This ability allows nuanced adaptation for individual client needs due to heterogeneity.
@@
To achieve the above, we propose “pMixFed”, a layer-wise, dynamic PFL approach that integrates Mixup zhang2017mixup between the shared global and personalized local models’ layers during both the broadcasting (global model sharing with local clients) and aggregation (aggregating distributed local models to update the global model) stages within a partial PFL framework. Our main contributions include:
  • We develop an online, dynamic interpolation method between local and global models using Mixup yoon2021fedmix, effectively addressing data heterogeneity and scalable across varying cohort sizes, degrees of data heterogeneity, and diverse model sizes and architectures.
@@
The remainder of this paper is organized as follows. In Section 2, we review the relevant literature on personalized federated learning, highlighting existing methodologies and their limitations. Section 3 introduces the problem formulation, including both fully personalized and partially personalized federated learning settings, accompanied by theoretical analysis. In Section 5, we present the proposed pMixFed framework, detailing the integration of adaptive and dynamic mixup strategies across different model layers. Section 6 provides extensive empirical evaluations of pMixFed, including comparisons with state-of-the-art baselines and analyses of computational and communication efficiency. Finally, Section 7 concludes the paper and outlines potential directions for future research.
@@ 2 Related Work
PFL seeks to adapt each client’s model to its individual data, preferences, and context. Similar to classic domain adaptation, data privacy can be an important concern in PFL stan2024unsupervised; verma2025leveraging; bingtao2025prodg; tian2024novel, the additional challenge, however, is that data is distributed and we need to adapt several models due to data heterogeneity between the clients. Since the advent of FL, a wide range of PFL methods have been proposed to address client heterogeneity (both statistical and system), which we group into four categories: (i) data-centric strategies; (ii) clustering-based approaches; (iii) global-model adaptation; and (iv) local-model personalization—the focus of this work, which also encompasses partial-model (parameter-decoupling) techniques. We discuss each category in turn below.
2.1 Data-Centric
Data-centric methods in PFL seek to mitigate distribution shift and class imbalance by shaping the data seen during training. for example, by adjusting sample sizes, label distributions, and client selection strategies tan2022towards. Common techniques include data normalization, feature engineering, augmentation, synthetic data generation, and adaptive client sampling. Representative examples include Astraea duan2019astraea, which performs reputation-based client selection by evaluating neighbors’ updates and rewarding trustworthy contributors; P2P k-SMOTE wang2021non, where clients synthesize minority-class examples via $k$-nearest-neighbor interpolation to reduce local imbalance; and FedMCCS 9212434, which improves fairness by selecting participants based on multiple factors (e.g., dataset size, class diversity, and training loss) rather than at random. FedAug jeong2018communication; zhao2018federated employs server-side data augmentation to alleviate heterogeneity, while FedHome wu2020fedhome augments local data using autoencoders trained with server-provided synthetic samples to address scarcity and imbalance. More recent works directly target class imbalance in PFL: Lee et al. lee2024regularizing construct a separate global model per class and let clients personalize by composing these class-wise globals according to local distributions; Liang et al. liang2024fedrema (FedReMa) selectively aggregates updates from the most relevant clients to enhance personalization under skewed data; and FedGA xiao2024fedga aligns client gradients before aggregation to reduce imbalance-induced bias and improve minority-class performance.
Despite their promise, data-centric approaches alter the underlying statistics of federated data and may inadvertently introduce bias or erase rare but important patterns. Several methods also rely on proxy or auxiliary server-side data, raising privacy and distribution-mismatch concerns. Finally, the added computation (and sometimes communication) can be burdensome for resource-constrained devices, limiting the practicality of these strategies in real-world PFL deployments.
2.2 Clustering
Clustering-based methods in PFL address the core challenge of heterogeneity: each client may hold data from markedly different distributions (e.g., distinct users, sensors, or hospitals). Naïvely averaging updates (e.g., FedAvg) can induce negative transfer when distributions diverge. Instead of a single global model, these approaches exploit structure across clients so that those with similar distributions reinforce one another while avoiding distortion from dissimilar updates. Clients thus retain the benefits of collaboration while mitigating negative transfer; in large populations, clustering/graph constructions also reduce complexity relative to fully personalized models. Clustering methods can be grouped into three categories: (1) Dynamic Clustering and Graph-based Aggregation; (2) Per-Client Weighting; and (3) Personalized Clustering.
In the first category, methods dynamically group clients or build graphs to personalize aggregation based on inter-client relationships. For example, FEDCEDAR wang2024rethinking builds dynamic weighted graphs among clients, enabling precise, personalized model distribution via dynamic graph propagation. Multi-Center FL long2023multi learns multiple global model “centers,” assigning each client to its best-fitting center through an EM-style procedure, thereby improving personalization under heterogeneous distributions. More recently, song2025chpfl proposes a three-layer clustered hierarchical PFL framework that adapts clustering at multiple levels to handle severe non-IID data.
In the second category (Per-Client Weighting), instead of uniform averaging, methods assign personalized aggregation weights according to client similarity or utility. FedSPD lin2024fedspd uses soft clustering so that clients converge toward cluster-relevant models while preserving personalization through flexible model consensus. FedDWA liu2023feddwa employs a parameter server to compute dynamic, per-client aggregation weights from model-update similarity, reducing communication and privacy overhead.
In the third category (Personalized Clustering), personalization is organized around classifier outputs or by merging multiple personalized models. For instance, pFedCk zhang2024personalized combines clustering with knowledge distillation: each client keeps a personalized model locally while an interaction model is shared on the server; features and logits are distilled across similar clients (clustered by the server) to enhance personalization. Likewise, COMET cho2023communication clusters clients based on distribution similarity and enables knowledge transfer through logits-based co-distillation, yielding cluster-specific personalized models while maintaining model heterogeneity and communication efficiency.
Clustering can be computationally heavy, and assignments may be unstable—clients can be misassigned when data are noisy or highly imbalanced. Maintaining multiple cluster models becomes costly at scale. Finally, similarity measures or EM responsibilities may leak distribution information, especially in centroid-based or hierarchical clustering.
2.3 Global Model Adaptation:
In these approaches, a single global model is maintained on the server and subsequently adapted to individual local models in a later phase. The primary goals are: (1) learning a robust, generalized model and (2) enabling fast, efficient local adaptation. Several techniques employ regularization terms to mitigate client drift and prevent model divergence during local updates. For example, FedProx li2020federated adds a proximal term so local updates stay close to the current global; SCAFFOLD karimireddy2020scaffold uses control variates to cancel client drift while training a single server model; MOON li2021model introduces a contrastive term to keep local representations aligned with the global and/or the previous local model; and FedCurv shoham2019overcoming infuses an EWC-style curvature penalty into the objective to keep local steps near a shared optimum.
Another line of single-global PFL approaches leverages meta-learning and transfer learning. Methods such as PerFedAvg fallah2020personalized learn a global initialization that each client can adapt with a few steps (MAML in FL). In FedL2P lee2023fedl2p meta-nets learn client-specific personalization hyperparameters (e.g., layerwise LRs / BN) to fine-tune the same global efficiently. MetaVers lim2024metavers meta-learns a versatile representation so the shared model adapts rapidly at clients. FedSteg yang2020fedsteg uses federated transfer/knowledge-transfer ideas for image steganalysis, and DPFed yu2020salvaging explicitly studies fine-tuning/local adaptation of a federated global model.
@@ 2.3 Global Model Adaptation:
Global model adaptation introduces specific challenges. Meta-learning methods can be computationally intensive, while regularization-based techniques add overhead by incorporating extra terms into the objective. Similarly, transfer-learning approaches can be communication-inefficient and often require a public dataset to enhance the server-side global model. A common limitation across most adaptation techniques is the need for a uniform model architecture across all clients, forcing devices with varying computational capabilities to use the same model size.
Another relevant direction is zero-shot learning (ZSL), which aims to train a model that generalizes on unseen classes, tasks, or domains without having direct training data for them hao2021towards; zhang2022fedzkt; rostami2022zero; asif2024advanced; guo2025feze. In the context of federated learning, ZSL can be interpreted as enabling the global or personalized models to perform well for new clients or unseen data distributions without requiring additional local training rounds. Recent works have leveraged generative models asif2024advanced to bridge the gap between seen and unseen tasks. Integrating ZSL capability into PFL could improve support for cold-start clients by allowing them to benefit from global knowledge without significant local updates.
2.4 Local Model Personalization:
The limitations of existing PFL methods have led to approaches that train customized models for each client. One line of work utilizes multi-task learning (MTL), a collaborative framework that facilitates information exchange across distinct tasks. for example, MOCHA smith2017federated, a canonical federated MTL method that trains related but distinct client models with a shared regularizer; and FedAMP huang2021personalized, where attentive message passing builds client-specific models with pairwise collaboration. In Ditto li2021ditto, each client optimizes its own personalized model with a regularization term. FEDHCA2 lu2024fedhca2 learns relationships between heterogeneous clients to produce personalized but collaborative models. In FedRes agarwal2020federated, clients learn local residuals that add to a shared global model; the personalized local predictor (global + local) can be viewed through the lens of MTL or parameter decoupling which is fully discussed below. Another category leverages knowledge distillation (KD) to support personalization when client-specific training objectives differ. In FedMD li2019fedmd, clients keep heterogeneous local models and distill via public logits to learn personalized local models. In FedGen zhu2021data, a server-side generator helps clients personalize via KD and feature transfer. FedGKT he2020group provides group teacher–student training and supports client-side students to form personalized models. These KD-based methods often include an adaptive fusion step that balances generalization and personalization without requiring full model sharing. Some approaches also combine KD with other categories; for example, pFedCK zhang2024personalized combines client clustering and KD to produce personalized client models, and FedGMKD zhang2024fedgmkd uses prototype-based KD with client-specific distillation to target local personalization.
A third, fast-growing line is hypernetwork or weight-generation–based personalization. Although the number of works is still relatively small, the space is expanding with adapters, prompts, and personalized parameter generation. The first paper, pFedHN shamsian2021personalized, trains a server-side hypernetwork that generates client-specific parameters from client embeddings; each client then receives its own weights, and personalization requires minimal client training. HyperFLoRA lu2024hyperflora trains a hypernetwork that generates LoRA adapter modules personalized to each client for LLM personalization: clients send lightweight statistics, and the server generates per-client adapters.
Both MTL- and KD-based approaches can incur significant computational and communication overhead, limiting scalability for large-scale FL deployments on resource-constrained devices. Many KD-based PFL methods assume access to proxy/public datasets for logit matching or generator training, which is unrealistic in privacy-sensitive settings. Hypernetworks, on the other hand, face a capacity bottleneck: a single hypernetwork must generate useful weights for many diverse clients; under high heterogeneity, it may underfit or collapse toward “average” solutions.
@@ 2.4 Local Model Personalization:
A prominent subcategory of local-model personalization, which decouples a shared parameter set (e.g., backbone) from client-specific modules (e.g., heads/adapters), is called partial models. This design is widely adopted because it is parameter-efficient, easy to quantize or deploy, and operationally transparent. By keeping the shared component stable and personalizing only a small subset, these methods reduce negative transfer across clients and mitigate catastrophic forgetting of global knowledge while enabling fast on-device adaptation. FedPer arivazhagan2019federated introduced partial models in FL, sharing only initial layers with generalized information and reserving final layers for personalization. FedBABU oh2021fedbabu divides the network into a shared body and a frozen head with fully connected layers. Other frameworks, such as FURL bui2019federated and LG-FedAvg liang2020think, apply partial PFL by retaining private feature embeddings or using compact representation learning for high-level features, respectively. Two baseline methods in this paper are FedAlt singhal2021federated and FedSim pillutla2022federated. FedAlt uses a stateless FL paradigm, reconstructing local models from the global model, while FedSim synchronously updates shared and local models with each iteration. FedSelect tamirisa2024fedselect is a Lottery-ticket-inspired subnetwork selection method that has been published recently. This method gradually personalizes a subnetwork per client while aggregating the rest. FedTSDP zhu2024federated is a PFL method that combines partial-model with clustering to better handle non-IID data. FDLoRA qi2024fdlora, which is an LLM-based FL systems, use LoRA (Low-Rank Adaptation) technique to improve personalization. FDLoRA introduces dual LoRA modules per client, one capturing global knowledge aggregated by the server, and the other capturing client-specific personalization. An adaptive fusion process reconciles the two, effectively balancing generalization with personalization without requiring full model sharing.
@@
These methods face challenges such as model update discrepancies (as shown in Figure 1) and catastrophic forgetting, where shared layers may undergo significant changes after only a few rounds of local training, resulting in sudden accuracy drops especially in high-scale Fl setting. Additionally, users with high personalization accuracy may freeze more layers than cold-start users or unreliable participants, who should rely more heavily on the global model. These challenges have motivated our development of a dynamic, adaptive layer-wise approach to balance generalization and personalization across clients, allowing tuning at different communication rounds to accommodate varying performance conditions.